### v2

@@
Ownership Chains in Multinational Enterprises thanks: We would like to thank Lionel Fontagné, Thierry Meyer, Ariell Reshef, and Farid Toubal for useful comments. We are also grateful to participants at the Royal Economic Society Annual Conference (Royal Economic Society) in 2024, the ASSA Annual Meeting in 2024, the European Trade Study Group (ETSG) in 2023, the Annual Conference on the Economics of Global Interactions in 2023, the Italian Trade Study Group in 2023, and the European Trade Study Group in 2022. Armando Rungi claims financial support from the PRIN project 2022W2245M, financed by the European Union - Next Generation EU.
Stefania Miricola Mail to: stefania.miricola@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems, IMT School for Advanced Studies, piazza San Francesco 19 - 55100 Lucca, Italy.    Armando Rungi Mail to: armando.rungi@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems, IMT School for Advanced Studies, piazza San Francesco 19 - 55100 Lucca, Italy.    Gianluca Santoni Mail to: gianluca.santoni@cepii.fr Centre d’Etudes Prospectives et d’Informations Internationales - Paris, France.
@@ 1 Introduction
To illustrate the structure of ownership chains visually, we present the case of a medium-sized MNE in Figure 1. QPS Holding has a corporate structure that spans three continents: the Americas, Asia, and Europe. QPS Holding is an American multinational active in bioanalytical, preclinical, and clinical research services. The parent company is at the top of the ownership hierarchy, with direct or indirect ownership of subsidiaries through four levels of hierarchy. For example, JSW Research DOO Beograd, located in Serbia, is owned by a chain that includes three intermediate subsidiaries: XDD ACQUISITION B.V., QPS Austria GMBH and QPS Hrvatska D.O.O.. This case illustrates the multi-tiered organization of multinational corporations that we describe throughout the paper, and the role of intermediate subsidiaries in cross-border corporate structures.
@@
  • Note: The figure shows the corporate structure of an MNE extracted from our sample. The node on top represents the parent company, other downstream nodes represent subsidiaries, and edges represent ownership linkages directed from one company to another. The parent company in the USA coordinates a chain across the Atlantic Ocean with an evident orientation from West to East, while directly owned subsidiaries on the other side of the Pacific Ocean are located in Taiwan.
@@ 2 Literature review
One of the fundamental explanations for hierarchical ownership structures is knowledge transmission and management efficiency. Garicano (2000) conceptualize firms as knowledge-based hierarchies in which managers’ hierarchical differentiation optimizes knowledge allocation. Routine problems are handled at lower levels, while complex problems are escalated. This framework has been extended to multinational firms, emphasizing the importance of structuring ownership networks to facilitate knowledge flows across geographically dispersed units (Antràs et al., 2006; Caliendo and Rossi-Hansberg, 2012). Gumpert (2018) show that firms face heterogeneous communication costs when transferring knowledge from headquarters to foreign affiliates, reinforcing the need for hierarchical complexity.  [^2] Other authors have used a similar theoretical framework. Altomonte et al. (2021) propose and test a model in which problem-solving efficiency helps determine the optimal hierarchical form of a corporate group, with parent firms supervising subsidiaries and managing communication among them. Altomonte and Rungi (2013) examine the relationship between hierarchical complexity and vertical integration and find consistencies with predictions from knowledge-based models. However, knowledge transfer alone does not fully explain ownership complexity. Beyond efficiency gains, firms must also retain control over dispersed operations, especially when facing institutional and contractual frictions. Gumpert et al. (2022) show that when the CEO’s ability to oversee operations is limited, firms introduce intermediate layers of management to optimize control. Similarly, Aghion and Tirole (1997) highlight the trade-off between corporate control and decentralized decision-making, shaping governance structures within firms.
@@
Indeed, governance concerns are central to multinational ownership structures. Firms must decide how much control to exercise over foreign affiliates. They must balance internal oversight with external contractual relationships. This decision is even more important when communication frictions and institutional weaknesses are taken into account. Giroud (2013) show that headquarters prefer investing in subsidiaries that are geographically closer due to lower monitoring costs, which in turn enhances productivity. Chauvin et al. (2024) reveal that time zone misalignment disrupts real-time coordination, increasing intra-firm communication costs. The role of contractual incompleteness is equally important. Ottaviano and Turrini (2007) show that weak contract enforcement significantly influences firms’ decisions on whether to own foreign affiliates or rely on arm’s-length relationships. When contracts are incomplete, MNEs prefer direct ownership over outsourcing to reduce hold-up risks, leading to more hierarchical structures. This perspective extends previous governance theories by showing that firms internalize operations not only to improve knowledge flows but also to mitigate legal uncertainties and transaction risks. Additionally, Bloom et al. (2012) provide empirical evidence that firms with high internal trust are more likely to delegate decision-making, which affects hierarchical organization.
@@
While corporations use hierarchical structures to maintain control and mitigate contractual risks, ownership chains are also shaped by regulatory considerations. The ability to shift profits across jurisdictions with different corporate tax rates provides a strong incentive for multinationals to create multi-layered ownership networks that facilitate the tax-efficient allocation of income. Unlike knowledge spillovers, which primarily improve internal efficiency, or control concerns, which focus on monitoring and governance, tax-motivated ownership complexity minimizes regulatory costs rather than operational coordination. Grubert (2003), highlighting the role of tax-motivated affiliate networks, show that a substantial amount of income shifting by U.S. MNEs occurs through intermediate holding companies that facilitate the routing of profits through conduit entities in tax-advantaged jurisdictions. More recently, Ferrari et al. (2022) develop a general equilibrium framework to analyze the impact of international tax rate differentials on corporate structuring. Their results show that the elasticity of shifted profits to tax rates is nearly three times larger than that of the tax base, underscoring how MNEs actively design ownership networks to minimize tax burdens rather than merely responding to production or operational needs. Empirical evidence confirms the large-scale impact of tax optimization on MNEs’ structuring. Laffitte and Toubal (2022) documents how firms strategically shift both sales and profits to tax havens, often disproportionately reporting revenues in low-tax jurisdictions while real economic activity remains elsewhere. Although tax optimization is a central driver of ownership complexity, other factors also play a role. [^3] For example, Lewellen and Robinson (2013) and Dyreng et al. (2015) show that U.S. multinationals often establish intermediate holding companies to facilitate regulatory arbitrage. Recent tax policy efforts, including the OECD’s Global Minimum Tax initiative, aim to reduce tax incentives for complex ownership networks by limiting tax rate differentials across jurisdictions (Ferrari et al., 2022). As global tax enforcement evolves, understanding the interplay between regulatory frameworks, tax planning and deal structuring remains critical.
@@
A key driver of multinational ownership is the need for hierarchical coordination of production and supply chains. Global companies structure their ownership networks to accommodate cross-jurisdictional interdependencies and ensure operational efficiency while maintaining management control over dispersed operations. Tintelnot (2017) develops a general equilibrium model showing that MNEs optimize production networks by trade-offs between the fixed costs of maintaining multiple locations and the efficiency gains from consolidating operations. Similarly, Arkolakis et al. (2018) and Wang (2021) show that MNEs make interdependent decisions about plant location, managerial control, and supply chain structure that are driven by economies of scale and coordination efficiencies. To facilitate technology transfer, sourcing, and logistics, firms often rely on intermediary subsidiaries that help bridge geographic and institutional gaps. Head and Mayer (2019) highlights how MNEs design supply chains to ensure brand consistency across markets, thereby reinforcing hierarchical control. These structures enhance continuity and efficiency in global value chains by reducing transaction costs associated with contractual incompleteness, supplier risk, and regulatory barriers. [^4] A related framework by Keller and Yeaple (2013) models firms’ decisions to embed technological knowledge in traded intermediates or to transfer it directly to foreign affiliates.
@@ 3 Mapping multinational ownership networks
When we refer to ownership structures, we mean the hierarchical control perimeter established by a parent company over its affiliates, which may be organized in ownership chains. The hypothetical chain of ownership in Figure 2 illustrates this concept. A subsidiary is directly controlled if the parent company holds an absolute majority of the voting rights in its shareholders’ meeting (e.g. parent company directly controls subsidiary A in Figure 2). Indirect control occurs when a subsidiary is owned through an intermediate affiliate that is itself controlled by the parent (e.g., Parent indirectly controls Subsidiaries B and C through an intermediary in Figure 2).
@@
Table 1: Geographic distribution of MNEs by host economy in 2019: Parent companies, intermediate affiliates, and final subsidiaries.
@@ 3.1 Stylized facts I: heterogeneity and relevance
  • Most subsidiaries by multinational enterprises are indirectly controlled through ownership chains.
@@
Complex multinational enterprises (MNEs) that operate through ownership chains account for a disproportionate share of global economic activity, accounting for 95% of total MNE revenues, despite representing only 27% of the firms in our dataset. In Figure 3, we distinguish between MNEs that exclusively use direct control over subsidiaries (left panel) and MNEs that include at least one ownership chain (right panel). While the latter group is numerically smaller, it dominates global sales, reinforcing the idea that larger corporate networks tend to be structured through multi-tiered ownership chains. [^7] Our findings are consistent with previous studies by Rungi et al. (2017); Altomonte and Rungi (2013), which show that the largest percentile of MNEs by corporate perimeter accounts for approximately 75% of global sales. This pattern emerges because MNEs without ownership chains tend to be smaller in size. On average, they generate 0.107 billion USD in revenue, compared with 1.782 billion USD for complex MNEs with chains of ownership. [^8] Ownership data tend to be more comprehensive than financial accounts because tax authorities require full disclosure of shareholding structures, while financial reporting requirements vary across jurisdictions. As a result, financial data may be incomplete due to exemptions and reporting discrepancies. In our analysis, we assume that missing sales data are not systematically correlated with firm size. Alternative imputation methods do not significantly alter our results, as confirmed in Table 3. Appendix Figure C1 illustrates that sample selection bias is often driven by geographic coverage rather than firm size.
@@
  • Note: Number of controlled subsidiaries per MNE on the x-axis and the frequency of MNEs on the y-axis. The first bar indicates the simplest MNEs made by one parent and one subsidiary. Complexity of ownership structure is possible only with more than two subsidiaries.
Table 3 provides further evidence of the heterogeneity of MNEs’ corporate structures. Looking first at the number of subsidiaries controlled, we find an average of 7.93 subsidiaries per parent company. But this masks a highly skewed distribution. The median MNE controls only one subsidiary, while the third quartile of MNEs reaches three subsidiaries per parent company. Moving to the right tail, we find that MNEs at the 90th percentile have more than 10 subsidiaries, and only at the 99th percentile do we reach the most complex MNEs, which control more than 120 subsidiaries, with a maximum of 5,075 subsidiaries worldwide. [^9] This is the case of Berkshire Hathaway, the largest multinational conglomerate holding company in the United States, ultimately owned by Warren Buffett.  [^10] In Appendix Figure C2, we estimate the distribution of controlled subsidiaries using a negative binomial regression, which yields a mean of 1.882 and an overdispersion parameter of 0.581. The figure visually represents a log-log plot of the frequency distribution of subsidiaries per parent. A similarly skewed pattern is observed when focusing on intermediate affiliates. The median MNE has only one intermediate affiliate, but the right tail expands sharply, with a jump from 2 to 546 subsidiaries between the 90th and 99th percentiles.
Looking at geographic dispersion, the table 3 shows that the average MNE operates in 2.81 countries. Since our definition of an MNE requires at least one foreign subsidiary, the minimum number of countries is two. However, the 90th percentile includes firms operating in four countries, while at the 99th percentile the number jumps to 19 countries and reaches a maximum of 147.111111This extreme case corresponds to DHL, the globally recognized logistics and courier group. Finally, industry diversification follows a similar pattern. The most complex MNEs, those in the right tail of the distribution, operate in more than 19 industries at the 99th percentile, reaching a maximum of 64 industries. [^12] According to our data, this is the case of HSBC Holdings PLC, originally founded as the Hongkong and Shanghai Banking Corporation and now a British banking group headquartered in London.

### v3

@@
Ownership Chains in Multinational Enterprises thanks: We would like to thank Lionel Fontagné, Thierry Mayer, Ariell Reshef, and Farid Toubal for useful comments. We are also grateful to participants at the Royal Economic Society Annual Conference (Royal Economic Society) in 2024, the ASSA Annual Meeting in 2024, the European Trade Study Group (ETSG) in 2023, the Annual Conference on the Economics of Global Interactions in 2023, the Italian Trade Study Group in 2023, and the European Trade Study Group in 2022. Armando Rungi claims financial support from the PRIN project 2022W2245M, financed by the European Union - Next Generation EU.
Stefania Miricola Mail to: stefania.miricola@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems, IMT School for Advanced Studies, piazza San Francesco 19 - 55100 Lucca, Italy.    Armando Rungi Mail to: armando.rungi@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems, IMT School for Advanced Studies, piazza San Francesco 19 - 55100 Lucca, Italy.    Gianluca Santoni Mail to: gianluca.santoni@cepii.fr. Centre d’Etudes Prospectives et d’Informations Internationales - Paris, France.
@@ 1 Introduction
To illustrate the structure of ownership chains visually, we present the case of a medium-sized MNE in Figure 1. QPS Holding has a corporate structure that spans three continents: the Americas, Asia, and Europe. QPS Holding is an American multinational active in bioanalytical, preclinical, and clinical research services. The parent company is at the top of the ownership hierarchy, with direct or indirect ownership of subsidiaries through four levels of hierarchy. For example, JSW Research DOO Beograd, located in Serbia, is owned by a chain that includes three intermediate subsidiaries: XDD Acquisition B.V., QPS Austria GMBH and QPS Hrvatska DOO. This case illustrates the multi-tiered organization of multinational corporations that we describe throughout the paper, and the role of intermediate subsidiaries in cross-border corporate structures.
@@
  • Note: The figure shows the corporate structure of an MNE extracted from our sample. The node on top represents the parent company, other downstream nodes represent subsidiaries, and arrows represent ownership linkages directed from one company to another. The parent company in the USA coordinates a chain across the Atlantic Ocean with an evident orientation from West to East, while directly owned subsidiaries on the other side of the Pacific Ocean are located in Taiwan.
@@ 2 Literature review
One of the fundamental explanations for the emergence of hierarchical structures is knowledge transmission and management efficiency. Garicano (2000) conceptualizes firms as knowledge-based hierarchies in which managers’ hierarchical differentiation optimizes knowledge allocation. Routine problems are handled at lower levels, while complex problems are escalated. This framework has been extended to multinational firms, emphasizing the importance of structuring the organization to facilitate knowledge flows across geographically dispersed units (Antràs et al., 2006; Caliendo and Rossi-Hansberg, 2012). Gumpert (2018) shows that firms face heterogeneous communication costs when transferring knowledge from headquarters to foreign affiliates, reinforcing the need for hierarchical complexity.222 Other authors have used a similar theoretical framework. Altomonte et al. (2021) propose and test a model in which problem-solving efficiency helps determine the optimal hierarchical form of a corporate group, with parent firms supervising subsidiaries and managing communication among them. Altomonte and Rungi (2013) examine the relationship between hierarchical complexity and vertical integration and find consistencies with predictions from knowledge-based models. However, knowledge transfer alone does not fully explain organizational complexity. Beyond efficiency gains, firms must also retain control over dispersed operations, especially when facing institutional and contractual frictions. Gumpert et al. (2022) show that when the CEO’s ability to oversee operations is limited, firms introduce intermediate layers of management to optimize control. Similarly, Aghion and Tirole (1997) highlight the trade-off between corporate control and decentralized decision-making, shaping governance structures within firms. Management literature has also emphasized that intermediate subsidiaries can serve as conduits for managerial coordination and knowledge flows within MNEs. Subsidiaries often act as critical nodes for internal knowledge transfer and coordination (Asmussen et al., 2011; Gupta and Govindarajan, 2000). Andersson et al. (2002) further highlight the strategic role of subsidiaries in developing capabilities and facilitating cross-unit learning within the group.
@@
Indeed, governance concerns extend to the case of global ownership structures. Aminadav and Papaioannou (2020) examine patterns of corporate control around the world, focusing on ultimate controlling shareholders, and show that control is strongly correlated with shareholder protection, the stringency of labor contracts, and the power of unions. Fonseca et al. (2023) show that baseline gravity models are effective in predicting bilateral country linkages, with more populous, wealthy, and proximate countries more likely to have control relationships. They also highlight the role of shared legal traditions, cultural ties, and historical connections. Our paper builds on existing work by examining the internal mechanisms of corporate control, in particular how control is exercised within multinational firms through ownership chains, the role of intermediate affiliates, and the challenges posed by communication frictions and geographic distance. In practice, parent firms must decide how much control to exert over their foreign affiliates. They must balance internal monitoring with external contractual relationships. This decision is even more important when communication frictions and institutional weaknesses are taken into account. Giroud (2013) shows that headquarters prefer investing in subsidiaries that are geographically closer due to lower monitoring costs, which in turn enhances productivity. Chauvin et al. (2024) reveal that time-zone misalignment disrupts real-time coordination, increasing intra-firm communication costs. The role of contractual incompleteness is equally important. Ottaviano and Turrini (2007) show that weak contract enforcement significantly influences firms’ decisions on whether to own foreign affiliates or rely on arm’s-length relationships. When contracts are incomplete, MNEs prefer direct ownership over outsourcing to reduce hold-up risks, leading to more hierarchical structures. This perspective extends previous governance theories by showing that firms internalize operations not only to improve knowledge flows but also to mitigate legal uncertainties and transaction risks. Additionally, Bloom et al. (2012) provide empirical evidence that firms with high internal trust are more likely to delegate decision-making, which affects hierarchical organization.
@@
While corporations use hierarchical structures to maintain control and mitigate contractual risks, ownership chains are also shaped by regulatory considerations. The ability to shift profits across jurisdictions with different corporate tax rates provides a strong incentive for multinationals to create multi-layered ownership networks that facilitate the tax-efficient allocation of income. Unlike knowledge spillovers, which primarily improve internal efficiency, or control concerns, which focus on monitoring and governance, tax-motivated ownership complexity minimizes regulatory costs rather than operational coordination. Grubert (2003), highlighting the role of tax-motivated affiliate networks, shows that a substantial amount of income shifting by US MNEs occurs through intermediate holding companies that facilitate the routing of profits through conduit entities in tax-advantaged jurisdictions. More recently, Ferrari et al. (2022) develop a general equilibrium framework to analyze the impact of international tax rate differentials on corporate structuring. Their results show that the elasticity of shifted profits to tax rates is nearly three times larger than that of the tax base, underscoring how MNEs actively design ownership networks to minimize tax burdens rather than merely responding to production or operational needs. Empirical evidence confirms the large-scale impact of tax optimization on MNEs’ structuring. Laffitte and Toubal (2022) document how firms strategically shift both sales and profits to tax havens, often disproportionately reporting revenues in low-tax jurisdictions while real economic activity remains elsewhere. Although tax optimization is a central driver of ownership complexity, other factors also play a role. [^3] For example, Lewellen and Robinson (2013) and Dyreng et al. (2015) show that US multinationals often establish intermediate holding companies to facilitate regulatory arbitrage. See also Francois and Vicard (2023) for recent evidence that complex ownership chains are associated with treaty shopping and tax-efficient investment routing. Recent tax policy efforts, including the OECD’s Global Minimum Tax initiative, aim to reduce tax incentives for complex ownership networks by limiting tax rate differentials across jurisdictions (Ferrari et al., 2022). As global tax enforcement evolves, understanding the interplay between regulatory frameworks, tax planning, and deal structuring remains critical.
@@
A key driver of multinational ownership is the need for hierarchical coordination of production and supply chains. Global companies structure their ownership networks to accommodate cross-jurisdictional interdependencies and ensure operational efficiency while maintaining management control over dispersed operations. Tintelnot (2017) develops a general equilibrium model showing that MNEs optimize production networks through trade-offs between the fixed costs of maintaining multiple locations and the efficiency gains from consolidating operations. Similarly, Arkolakis et al. (2018) and Wang (2021) show that MNEs make interdependent decisions about plant location, managerial control, and supply chain structure that are driven by economies of scale and coordination efficiencies. To facilitate technology transfer, sourcing, and logistics, firms often rely on intermediary subsidiaries that help bridge geographic and institutional gaps. Head and Mayer (2019) highlight how MNEs design supply chains to ensure brand consistency across markets, thereby reinforcing hierarchical control. These structures enhance continuity and efficiency in global value chains by reducing transaction costs associated with contractual incompleteness, supplier risk, and regulatory barriers. [^4] A related framework by Keller and Yeaple (2013) models firms’ decisions to embed technological knowledge in traded intermediates or to transfer it directly to foreign affiliates.
@@ 3 Mapping multinational ownership networks
When we refer to ownership structures, we mean the hierarchical control perimeter established by a parent company over its affiliates, which may be organized in ownership chains. The hypothetical chain of ownership in Figure 2 illustrates this concept. A subsidiary is directly controlled if the parent company holds an absolute majority of the voting rights in its shareholders’ meeting (e.g. parent company directly controls subsidiary A in Figure 2). Indirect control occurs when a subsidiary is owned through an intermediate affiliate that is itself controlled by the parent (e.g. Parent indirectly controls Subsidiaries B and C through an intermediary in Figure 2).
@@
Table 1: Geographic distribution of MNEs by host economy in 2019: Parent companies, intermediate affiliates, and final subsidiaries
@@ 3.1 Stylized facts I: heterogeneity and relevance
  • Most subsidiaries of multinational enterprises are indirectly controlled through ownership chains.
@@
Complex multinational enterprises that operate through ownership chains account for a disproportionate share of global economic activity, accounting for 95% of total MNE revenues, despite representing only 27% of the firms in our dataset. In Figure 3, we distinguish between MNEs that exclusively use direct control over subsidiaries (left panel) and MNEs that include at least one ownership chain (right panel). While the latter group is numerically smaller, it dominates global sales, reinforcing the idea that larger corporate networks tend to be structured through multi-tiered ownership chains. [^7] Our findings are consistent with previous studies by Rungi et al. (2017); Altomonte and Rungi (2013), which show that the largest percentile of MNEs by corporate perimeter accounts for approximately 75% of global sales. This pattern emerges because MNEs without ownership chains tend to be smaller in size. On average, they generate 0.107 billion USD in revenue, compared with 1.782 billion USD for complex MNEs with chains of ownership. [^8] Ownership data tend to be more comprehensive than financial accounts because tax authorities require full disclosure of shareholding structures, while financial reporting requirements vary across jurisdictions. As a result, financial data may be incomplete due to exemptions and reporting discrepancies. In our analysis, we assume that missing sales data are not systematically correlated with firm size. Alternative imputation methods do not significantly alter our results, as confirmed in Table 3. Appendix Figure C1 illustrates that sample selection bias is often driven by geographic coverage rather than firm size.
@@
  • Note: Number of controlled subsidiaries per MNE on the x-axis and the frequency of MNEs on the y-axis. The first bar indicates the simplest MNEs, consisting of one parent and one subsidiary. Complexity of ownership structure is possible only with more than two subsidiaries.
Table 3 provides further evidence of the heterogeneity of MNEs’ corporate structures. Looking first at the number of subsidiaries controlled, we find an average of 7.93 subsidiaries per parent company. But this masks a highly skewed distribution. The median MNE controls only one subsidiary, while the third quartile of MNEs reaches three subsidiaries per parent company. Moving to the right tail, we find that MNEs at the 90th percentile have more than 10 subsidiaries, and only at the 99th percentile do we reach the most complex MNEs, which control more than 120 subsidiaries, with a maximum of 5,075 subsidiaries worldwide. [^9] This is the case of Berkshire Hathaway, the largest multinational conglomerate holding company in the United States, ultimately owned by Warren Buffett.  [^10] In Appendix Figure C2, we estimate the distribution of controlled subsidiaries using a negative binomial regression, which yields a mean of 1.882 and an overdispersion parameter of 0.581. The figure visually represents a log-log plot of the frequency distribution of subsidiaries per parent. A similarly skewed pattern is observed when focusing on intermediate affiliates. The median MNE has zero intermediate affiliate, but the right tail expands sharply, with a jump from 2 to 20 subsidiaries between the 90th and 99th percentiles.
Looking at geographic dispersion, Table 3 shows that the average MNE operates in 2.81 countries. Since our definition of an MNE requires at least one foreign subsidiary, the minimum number of countries is two. However, the 90th percentile includes firms operating in four countries, while at the 99th percentile the number jumps to 19 countries and reaches a maximum of 147.111111This extreme case corresponds to DHL, the globally recognized logistics and courier group. Finally, industry diversification follows a similar pattern. The most complex MNEs, those in the right tail of the distribution, operate in more than 15 industries at the 99th percentile, reaching a maximum of 64 industries. [^12] According to our data, this is the case of HSBC Holdings PLC, originally founded as the Hongkong and Shanghai Banking Corporation and now a British banking group headquartered in London.